# HexGen: Generative Inference of Large-Scale Foundation Model over Heterogeneous Decentralized Environment

### v2

@@ 4.3. Searching via Genetic Algorithm
Put it together. We adopt a standard iterative procedure of the genetic algorithm: for each iteration, we conduct the mutation operation(s), run the dynamic algorithm to determine the optimal pipeline assignments, and adjust the pipeline partitions for the current off-springs. To estimate the expected SLO, we adopt the inference task simulator from AlpaServe (Li et al., 2023). To deploy HexGen in a real heterogeneous decentralized environment, we implement a task coordinator. The task coordinator manages the GPUs from the heterogeneous pool and organizes the independent pipeline parallel worker groups according to the optimal allocation based on the approach we introduced in Section 4. Concretely, the task coordinator is mainly based on an open-source implementation of decentralized computation coordination (Yao, 2023) that utilizes libP2P (LibP2P, 2023) to establish connections among the work groups in a peer-to-peer network. When an inference request is received by the task coordinator, the request will be directed to an appropriate worker group according to the scheduling result.
5. Evaluation
Figure 2. SLO attainment results to evaluate cost performance trade-offs. Each row corresponds to a particular output sequence length (32, 64, 128). The first four columns correspond to different SLO scales ranging from 8 to 0.125 requests per second. The last column represents the performance comparison of various settings at different request rates.
@@ 5.1. Experimental Setup
Evaluation metrics. We evaluate system performance based on SLO attainment: we generate inference workload according to a Poisson process parameterized by the request rate. The consecutive requests (inter-arrival times) follow an exponential distribution. For a target SLO goal (e.g., $99\%$), we focus on two metrics: i) the minimum latency deadline required to achieve the desired attainment, and ii) the system’s resilience to peak request rate. We apply the most popular open-source Llama-2 (70B) model on some real-world prompts (Lmsys, 2023), and test output sequence lengths from $32$ to $128$, and request rates varying between $0.125$ $10$ requests per second.
5.2. Cost Performance Trade-off
@@
Cost efficiency. When given the relatively same budget, HexGen under the full budget in the heterogeneous setting clearly outperforms FlashAttention in the homogeneous datacenter setting. In fact, HexGen reaches up to 2.3×2.3$\times$2.3 × and on average 1.5×1.5$\times$1.5 × lower latency deadlines, and is capable of handling a peak request rate that is up to 4×4$\times$4 × higher (2×2$\times$2 × higher on average). Furthermore, even when we reduce the budget in the heterogeneous setting by half, HexGen still reveals similar performance to FlashAttention in the homogeneous setting. We believe that this is strong evidence to illustrate that a decentralized system such as HexGen is capable of managing heterogeneous GPUs to provide more economical foundation model inference services without compromising service quality.
@@ 5.3. End-to-end System Performance
Figure 3. HexGen and Petals. Two rows correspond to output sequence lengths of $32$ and $64$. First two columns illustrate the results of different SLO scales. The last column shows the effects of request rate on SLO attainment.
@@
Figure 4. HexGen v.s. Huggingface-TGI. Two rows represent output sequence lengths of $32$ and $64$. First two columns show the result of different SLO scales. The last column shows the effects of request rate on SLO attainment.
Compare with Huggingface-TGI. We also compare HexGen under the full budget in heterogeneous setting with Huggingface-TGI in the homogeneous data center settings. In this case, HexGen gets almost the same end-to-end performance in both latency and request handling, achieves up to 1.25×1.25$\times$1.25 × lower latency deadlines, and is able to handle requests at the same level of rates. When we compare HexGen with Huggingface-TGI, we realize that one particular reason that actually limits HexGen from reaching its full potential when compared with Huggingface-TGI — state-of-the-art foundation model inference services usually include some advanced batching policy to improve the hardware efficiency significantly; the current version of HexGen has not integrated this feature yet since the batching policy under the heterogeneous setting issues some unique challenges given the diversified executing time and memory limit of each independent pipeline groups, which makes the current batching policy difficult to integrate. We acknowledge this limitation in the current version of HexGen and leave this as an important future work to further improve HexGen’s end-to-end inference performance under the heterogeneous setting.
5.4. Effectiveness of the Scheduling Algorithm
@@
Figure 5. Convergence comparison of the proposed search strategy and random mutation.
@@
To evaluate the effectiveness of the scheduling algorithm, we disable the advanced mutation strategies introduced in Section 4.3 by replacing this part with random mutation and comparing the convergence behavior with the proposed algorithm. We benchmarked the full-price and half-price cases, where we set the output sequence length to $32$, and the SLO scale as $5$. Figure 5 illustrates the result. We see that our proposed constrained mutation policy significantly outperforms random mutation — the proposed strategy can find assignments that can manage around $26\%$ more SLO attainments than random mutation and converges much faster, while random mutation gets stuck in some local minimum. Additionally, we verified that in both cases, the estimated SLO attainments ($92\%$ and $82\%$) closely align with the actual attainments ($94\%$ and $86\%$) demonstrated in Figure 2.
6. Related Work
6.1. Foundation model inference optimization
@@
There have been many efforts to accelerate the inference service in terms of both system optimization and algorithm design. On the system side, research efforts have focused on enhancing hardware efficiency through meticulous system optimizations (Fang et al., 2021b; Yu et al., 2022; Li et al., 2023; Kwon et al., 2023; Dao et al., 2022). For example, AlpaServe (Li et al., 2023) proposes a concrete analysis of model parallel strategies and model placement to improve inference service efficiency; PagedAttention (Kwon et al., 2023) introduces an advanced memory management system to batch inference jobs inspired by the classic design of virtual memory and paging; FLashAttention (Dao, 2023) leveraged GPU memory hierarchy to reduce GPU memory access significantly, leading to runtime speedup. On the algorithm side, some advanced algorithm designs have also been proposed (Leviathan et al., 2023; Yao et al., 2022; Liu et al., 2023). For example, speculative decoding (Leviathan et al., 2023; Miao et al., 2023a; Spector and Re, 2023) based algorithms improve the system efficiency by leveraging a small approximation model for prediction and the original model for parallel verification. A similar idea has been explored by Medusa (Cai et al., 2023), which implements a multiple-head decoding mechanism for parallel token verification. Low precision computation has also been explored to speed up generative inference, such as quantization (Yao et al., 2022; Frantar et al., 2022; Xiao et al., 2022; Lin et al., 2023), sparsification (Frantar and Alistarh, 2023; Liu et al., 2023) and distillation (Kwon et al., 2022).
6.2. Decentralized computation platform
@@
Distributed machine learning becomes an important workflow for data management and distributed systems (Yuan et al., 2021; Fang et al., 2021a; Miao et al., 2021; Gan et al., 2021; Fu et al., 2022; Miao et al., 2022; Xie et al., 2023; Zhang et al., 2022). Recently, there have been some emerging research attempts on deploying machine learning computations across a variety of decentralized and heterogeneous computational resources (Ali et al., 2022; Miao et al., 2023b; Zhang et al., 2023). For example, sky computing (Stoica and Shenker, 2021; Yang et al., 2023) builds an additional layer above classic cloud platforms to access more economic computation resources and enable interoperability between multi-clouds; SAKSHI (Bhat et al., 2023) has proposed a blueprint to advocate for the hosting and delivery of reliable AI models in the landscape of AI services. There have also been some attempts to deploy machine learning training in a collaborative environment (Diskin et al., 2021; Yuan et al., 2022; Ryabinin et al., 2023). However, most of this work does not focus on the system implementation and scheduling for generative inference workflows. Perhaps the most relevant effort is Petals (Borzunov et al., 2022), which allows users to donate different GPUs to perform inference and small-scale fine-tuning. However, Petals is mainly based on dynamic coordination from swarm parallelism (Ryabinin et al., 2023), whose performance is limited by the lack of scheduling of the decentralized inference as we illustrated in Section 5.3.
7. Conclusion
@@
In this paper, we explore the opportunity to deploy the inference service of foundation models via a heterogeneous regime with devices of different computation capacities connected over a heterogeneous network. Toward this end, we propose HexGen, a generative inference framework with asymmetric parallel support and an effective scheduling algorithm to accommodate such deployment. Our empirical study suggests that when given the same budget, HexGen can outperform the centralized homogeneous deployment by either achieving 2.3×2.3$\times$2.3 × lower latency deadlines or tolerating up to 4×4$\times$4 × more traffic request rate; additionally, HexGen also significantly outperforms Petals, the state-of-the-art decentralized collaborative inference prototype by 10×10$\times$10 × more traffic request rate.
We will make the HexGen system fully open-sourced and hope that such a system can contribute to the democratization of the usage of foundation models.
@@ Appendix A Summarization of Notations
Table 2. Table of notations.
@@ Appendix B Modeling the Generative Inference Cost
(4) $\textsc{C}^{i,j}_{\text{comp}}\left(\mathbf{d}_{i,j}\right)=\max_{d\in\mathbf{d}_{i,j}}\left(\frac{12H^{2}B_{\text{type}}s_{t}^{\text{out}}}{\left|\mathbf{d}_{i,j}\right|m_{d}}\right)\cdot l_{i,j}+\max_{d\in\mathbf{d}_{i,j}}\left(\frac{24b_{t}\left(s^{\text{in}}_{t}+s^{\text{out}}_{t}\right)H^{2}}{\left|\mathbf{d}_{i,j}\right|c_{d}}\right)\cdot l_{i,j}$
@@
(5) $\textsc{C}^{i,j}_{\text{comm-tp}}\left(\mathbf{d}_{i,j}\right)=$ $\max_{d\in\mathbf{d}_{i,j}}\left(\sum_{d^{\prime}\in\mathbf{d}_{i,j}-\{d\}}\left(\alpha_{d,d^{\prime}}+\frac{b_{t}s^{\text{in}}_{t}HB_{\text{type}}}{{\left|\mathbf{d}_{i,j}\right|\beta}_{d,d^{\prime}}}\right)\right)\cdot 4l_{i,j}$
@@
(6) $\textsc{C}^{i,j}_{\text{comm-pp}}\left(\mathbf{d}_{i,j}\right)=\min_{d\in\mathbf{d}_{i,j},d^{\prime}\in\mathbf{d}_{i,j+1}}\left(\alpha_{d,d^{\prime}}+\frac{b_{t}s^{\text{in}}_{t}HB_{\text{type}}}{\beta_{d,d^{\prime}}}\right)+\min_{d\in\mathbf{d}_{i,j},d^{\prime}\in\mathbf{d}_{i,j+1}}\left(\alpha_{d,d^{\prime}}+\frac{b_{t}HB_{\text{type}}}{\beta_{d,d^{\prime}}}\right)\cdot s_{t}^{\text{out}}$
@@
(7) $\begin{split}\textsc{C}^{d}_{\text{mem}}\left(\mathbf{d}_{i,j}\right)=\left(\frac{12H^{2}B_{\text{type}}}{\left|\mathbf{d}_{i,j}\right|}+\frac{2b_{t}\left(s^{\text{in}}_{t}+s^{\text{out}}_{t}\right)HB_{\text{type}}}{\left|\mathbf{d}_{i,j}\right|}\right)\times l_{i,j}+4b_{t}\left(s^{\text{in}}_{t}+s^{\text{out}}_{t}\right)HB_{\text{type}}\end{split}$

### v3

@@ 4.3 Searching via Genetic Algorithm
Put it together. We adopt a standard iterative procedure of the genetic algorithm: for each iteration, we conduct the mutation operation(s), run the dynamic algorithm to determine the optimal pipeline assignments, and adjust the pipeline partitions for the current off-springs. To estimate the expected SLO, we adopt the inference task simulator from AlpaServe (Li et al., 2023).
5 Evaluation
Figure 2: SLO attainment results to evaluate cost performance trade-offs. Each row corresponds to a particular output sequence length (32, 64, 128). The first four columns correspond to different SLO scales ranging from 8 to 0.125 requests per second. The last column represents the performance comparison of various settings at different request rates.
@@ 5.1 Experimental Setup
Evaluation metrics. Following the generative inference evaluation setup from AlpaServe (Li et al., 2023), we test system performance based on SLO attainment, refers to the proportion of requests that can be finished within a predefined performance threshold. Specifically, we measure SLO attainment in terms of the percentage of requests served within the time frame set by the SLO. We scale the SLO to various multiples of the execution latency of A100 GPUs (SLO Scale in Figure 2), which allows us to evaluate system performance under different levels of operational stringency. We generate inference workload according to a Poisson process parameterized by the request rate. The consecutive requests (inter-arrival times) follow an exponential distribution. For a target SLO goal (e.g., $99\%$), we focus on two metrics: i) the minimum latency deadline required to achieve the desired attainment, and ii) the system’s resilience to peak request rate. We apply the most popular open-source Llama-2 (70B) model on some real-world prompts (Lmsys, 2023), and test output sequence lengths from $32$ to $128$, and request rates varying between $0.125$ $10$ requests per second.
5.2 Cost Performance Trade-off
@@
Cost efficiency. When given the relatively same budget, HexGen under the full budget in the heterogeneous setting clearly outperforms FlashAttention in the homogeneous datacenter setting. In fact, HexGen reaches up to 2.3×2.3$\times$2.3 × and on average 1.5×1.5$\times$1.5 × lower latency deadlines, and is capable of handling a peak request rate that is up to 4×4$\times$4 × higher (2×2$\times$2 × higher on average). Specifically, by analyzing the scheduling results [^5] All the strategies chosen by the scheduling algorithm can be found in Appendix F. for heterogeneous-full-price cases, we find that our scheduling approach always prioritizes intra-machine tensor model parallelism to minimize single request latency and employs inter-machine pipeline parallelism to reduce communication over limited bandwidth. It avoids cross-region communication due to ultra-low bandwidth and aims to maximize device memory utilization by incorporating as many model replicas as possible, thereby enhancing parallel request processing. Furthermore, even when we reduce the budget in the heterogeneous setting by half, HexGen still reveals similar performance to FlashAttention in the homogeneous setting. We believe that this is strong evidence to illustrate that a decentralized system such as HexGen is capable of managing heterogeneous GPUs to provide more economical foundation model inference services without compromising service quality.
@@ 5.3 End-to-end System Performance
Figure 3: HexGen and Petals. Two rows correspond to output sequence lengths of $32$ and $64$. First two columns illustrate the results of different SLO scales. The last column shows the effects of request rate on SLO attainment.
@@
To evaluate HexGen over dynamic GPU pools, we test a scenario where 4 GPUs leave the current allocation scheduled by HexGen. In this case, HexGen will re-run the search algorithm to find the new optimal allocation. Interestingly, we find this simple policy is very effective—the genetic algorithm is based on local search, which demands much less iteration of searching when only a small portion of GPUs dynamically join or leave. We notice that HexGen can rerun the searching algorithm in less than 30 seconds to find the new optimal allocation. Figure 4 illustrates HexGen’s performance before and after the 4 GPUs become offline. We see the performance gap is considerably small under such dynamics. In addition, we find that the performance of HexGen with 4 GPUs offline is still significantly better than Petals.
Figure 4: SLO attainment results of HexGen compared with HexGen with 4 GPUs offline.
@@
Figure 5: HexGen v.s. Huggingface-TGI. Two rows represent output sequence lengths of $32$ and $64$. First two columns show the result of different SLO scales. The last column shows the effects of request rate on SLO attainment.
Compare with Huggingface-TGI. We also compare HexGen under the full budget in heterogeneous setting with Huggingface-TGI in the homogeneous data center settings. In this case, HexGen gets almost the same end-to-end performance in both latency and request handling, achieves up to 1.25×1.25$\times$1.25 × lower latency deadlines, and is able to handle requests at the same level of rates.
5.4 Effectiveness of the Scheduling Algorithm
@@
Figure 6: Convergence comparison of the proposed search strategy and random mutation.
@@
To evaluate the effectiveness of the scheduling algorithm, we disable the advanced mutation strategies introduced in Section 4.3 by replacing this part with random mutation and comparing the convergence behavior with the proposed algorithm. We benchmarked the full-price and half-price cases, where we set the output sequence length to $32$, and the SLO scale as $5$. Figure 6 illustrates the result. The proposed search strategy takes 2.1 and 1.5 minutes to identify the optimal assignments for full-price and half-price heterogeneous scenarios. This search process runs once before the system is initially deployed, which makes its time cost negligible. We see that our proposed constrained mutation policy significantly outperforms random mutation — the proposed strategy can find assignments that can manage around $26\%$ more SLO attainments than random mutation and converges much faster, while random mutation gets stuck in some local minimum. Additionally, we verified that in both cases, the estimated SLO attainments ($92\%$ and $82\%$) closely align with the actual attainments ($94\%$ and $86\%$) demonstrated in Figure 2. We also provide a direct performance comparison between random initialized allocation (the allocation after executing the initialization method mentioned in subsection 4.3), random mutated policy through evolution and HexGen’s result in the heterogeneous-half-price scenario. The results are illustrated in Figure 7.
Figure 7: SLO attainment results of HexGen compared with random mutation based results and results w/o any searching efforts.
6 Related Work
@@
Foundation model inference optimization. There have been many efforts to accelerate the inference service in terms of both system optimization and algorithm design. On the system side, research efforts have focused on enhancing hardware efficiency through meticulous system optimizations (Fang et al., 2021; Yu et al., 2022; Li et al., 2023; Kwon et al., 2023; Dao et al., 2022). On the algorithm side, some advanced algorithm designs have also been proposed (Leviathan et al., 2023; Yao et al., 2022; Liu et al., 2023), including speculative decoding (Spector & Re, 2023), multiple-head decoding mechanism (Cai et al., 2023) and low precision computation such as quantization (Yao et al., 2022; Frantar et al., 2022; Xiao et al., 2022; Lin et al., 2023), sparsification (Frantar & Alistarh, 2023; Liu et al., 2023) and distillation (Kwon et al., 2022).
@@
Decentralized computation platform. Recently, there have been some emerging research attempts on deploying machine learning computations across a variety of decentralized and heterogeneous computational resources (Ali et al., 2022; Miao et al., 2023b; Zhang et al., 2023; Stoica & Shenker, 2021; Bhat et al., 2023), e.g., distributed training in a collaborative environment (Diskin et al., 2021; Yuan et al., 2022; Ryabinin et al., 2023). However, most of this work does not focus on the system implementation and scheduling for generative inference workflows. Perhaps the most relevant effort is Petals (Borzunov et al., 2022), which allows users to donate different GPUs to perform inference and small-scale fine-tuning. However, Petals is mainly based on dynamic coordination from swarm parallelism (Ryabinin et al., 2023), whose performance is limited by the lack of scheduling of the decentralized inference.
7 Conclusion
@@
In this paper, we explore the opportunity to deploy the inference service of foundation models via a heterogeneous regime with devices of different computation capacities connected over a heterogeneous network. Toward this end, we propose HexGen, a generative inference framework with asymmetric parallel support and an effective scheduling algorithm to accommodate such deployment. Our empirical study suggests that when given the same budget, HexGen can outperform the centralized homogeneous deployment by either achieving 2.3×2.3$\times$2.3 × lower latency deadlines or tolerating up to 4×4$\times$4 × more traffic request rate; additionally, HexGen also significantly outperforms Petals, the state-of-the-art decentralized collaborative inference prototype by 10×10$\times$10 × more traffic request rate. We will make the HexGen system fully open-sourced and hope that such a system can contribute to the democratization of the usage of foundation models.
Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Appendix A Summarization of Notations
Table 2: Table of notations.
@@ Appendix B Modeling the Generative Inference Cost
$\textsc{C}^{i,j}_{\text{comp}}\left(\mathbf{d}_{i,j}\right)=\max_{d\in\mathbf{d}_{i,j}}\left(\frac{12H^{2}B_{\text{type}}s_{t}^{\text{out}}}{\left|\mathbf{d}_{i,j}\right|m_{d}}\right)\cdot l_{i,j}+\max_{d\in\mathbf{d}_{i,j}}\left(\frac{24b_{t}\left(s^{\text{in}}_{t}+s^{\text{out}}_{t}\right)H^{2}}{\left|\mathbf{d}_{i,j}\right|c_{d}}\right)\cdot l_{i,j}$ (4)
@@
$\textsc{C}^{i,j}_{\text{comm-tp}}\left(\mathbf{d}_{i,j}\right)=$ $\max_{d\in\mathbf{d}_{i,j}}\left(\sum_{d^{\prime}\in\mathbf{d}_{i,j}-\{d\}}\left(\alpha_{d,d^{\prime}}+\frac{b_{t}s^{\text{in}}_{t}HB_{\text{type}}}{{\left|\mathbf{d}_{i,j}\right|\beta}_{d,d^{\prime}}}\right)\right)\cdot 4l_{i,j}$ (5)
@@
$\textsc{C}^{i,j}_{\text{comm-pp}}\left(\mathbf{d}_{i,j}\right)=\min_{d\in\mathbf{d}_{i,j},d^{\prime}\in\mathbf{d}_{i,j+1}}\left(\alpha_{d,d^{\prime}}+\frac{b_{t}s^{\text{in}}_{t}HB_{\text{type}}}{\beta_{d,d^{\prime}}}\right)+\min_{d\in\mathbf{d}_{i,j},d^{\prime}\in\mathbf{d}_{i,j+1}}\left(\alpha_{d,d^{\prime}}+\frac{b_{t}HB_{\text{type}}}{\beta_{d,d^{\prime}}}\right)\cdot s_{t}^{\text{out}}$ (6)
@@
$\begin{split}\textsc{C}^{d}_{\text{mem}}\left(\mathbf{d}_{i,j}\right)=\left(\frac{12H^{2}B_{\text{type}}}{\left|\mathbf{d}_{i,j}\right|}+\frac{2b_{t}\left(s^{\text{in}}_{t}+s^{\text{out}}_{t}\right)HB_{\text{type}}}{\left|\mathbf{d}_{i,j}\right|}\right)\times l_{i,j}+4b_{t}\left(s^{\text{in}}_{t}+s^{\text{out}}_{t}\right)HB_{\text{type}}\end{split}$ (7)
Performance alignment. To evaluate the accuracy of our cost model, we conduct a suite of micro-benchmarks to validate the output of our cost estimation by comparing it with the actual execution time. The results are listed in Table 3 and suggest that our cost model can align with the actual execution accurately.
Table 3: Comparison of Benchmarked and Estimated Performance Metrics.
Input/Output
Length
Parallel
Configuration
Prefill Time -
Benchmarked
Prefill Time -
Estimated
Decode Time -
Benchmarked
Decode Time -
Estimated
256/32 TP=8 2.72s 2.99s 2.43s 2.46s
TP=4 PP=2 3.79s 3.85s 2.25s 2.14s
TP=2 PP=4 5.26s 5.25s 3.29s 3.04s
PP=8 8.04s 7.83s 6.04s 5.60s
512/64 TP=8 3.04s 3.10s 4.76s 4.92s
TP=4 PP=2 4.16s 4.10s 4.32s 4.28s
TP=2 PP=4 5.57s 5.63s 6.65s 6.08s
PP=8 8.27s 8.49s 12.4s 11.2s
Appendix C Task Coordinator
To deploy HexGen in a real heterogeneous decentralized environment, we implement a task coordinator. The task coordinator manages the GPUs from the heterogeneous pool and organizes the independent pipeline parallel worker groups according to the optimal allocation based on the approach we introduced in Section 4. Concretely, the task coordinator is mainly based on an open-source implementation of decentralized computation coordination (Yao, 2023) that utilizes libP2P (LibP2P, 2023) to establish connections among the work groups in a peer-to-peer network. When an inference request is received by the task coordinator, the request will be directed to an appropriate worker group according to the scheduling result.
Appendix D Limitation of Batching Implementation
When we compare HexGen with Huggingface-TGI, we realize that one particular reason that actually limits HexGen from reaching its full potential when compared with Huggingface-TGI — state-of-the-art foundation model inference services usually include some advanced batching policy to improve the hardware efficiency significantly; the current version of HexGen has not integrated this feature yet since the batching policy under the heterogeneous setting issues some unique challenges given the diversified executing time and memory limit of each independent pipeline groups, which makes the current batching policy difficult to integrate. We acknowledge this limitation in the current version of HexGen and leave this as an important future work to further improve HexGen’s end-to-end inference performance under the heterogeneous setting.
Appendix E Extended Related Works
Foundation model inference optimization. There have been many efforts to accelerate the inference service in terms of both system optimization and algorithm design. On the system side, research efforts have focused on enhancing hardware efficiency through meticulous system optimizations (Fang et al., 2021; Yu et al., 2022; Li et al., 2023; Kwon et al., 2023; Dao et al., 2022). For example, AlpaServe (Li et al., 2023) proposes a concrete analysis of model parallel strategies and model placement to improve inference service efficiency; PagedAttention (Kwon et al., 2023) introduces an advanced memory management system to batch inference jobs inspired by the classic design of virtual memory and paging; FLashAttention (Dao, 2023) leveraged GPU memory hierarchy to reduce GPU memory access significantly, leading to runtime speedup. On the algorithm side, some advanced algorithm designs have also been proposed (Leviathan et al., 2023; Yao et al., 2022; Liu et al., 2023). For example, speculative decoding (Leviathan et al., 2023; Miao et al., 2023a; Spector & Re, 2023) based algorithms improve the system efficiency by leveraging a small approximation model for prediction and the original model for parallel verification. A similar idea has been explored by Medusa (Cai et al., 2023), which implements a multiple-head decoding mechanism for parallel token verification. Low precision computation has also been explored to speed up generative inference, such as quantization (Yao et al., 2022; Frantar et al., 2022; Xiao et al., 2022; Lin et al., 2023), sparsification (Frantar & Alistarh, 2023; Liu et al., 2023) and distillation (Kwon et al., 2022).
Decentralized computation platform. Recently, there have been some emerging research attempts on deploying machine learning computations across a variety of decentralized and heterogeneous computational resources (Ali et al., 2022; Miao et al., 2023b; Zhang et al., 2023). For example, sky computing (Stoica & Shenker, 2021; Yang et al., 2023) builds an additional layer above classic cloud platforms to access more economic computation resources and enable interoperability between multi-clouds; SAKSHI (Bhat et al., 2023) has proposed a blueprint to advocate for the hosting and delivery of reliable AI models in the landscape of AI services. There have also been some attempts to deploy machine learning training in a collaborative environment (Diskin et al., 2021; Yuan et al., 2022; Ryabinin et al., 2023). However, most of this work does not focus on the system implementation and scheduling for generative inference workflows. Perhaps the most relevant effort is Petals (Borzunov et al., 2022), which allows users to donate different GPUs to perform inference and small-scale fine-tuning. However, Petals is mainly based on dynamic coordination from swarm parallelism (Ryabinin et al., 2023), whose performance is limited by the lack of scheduling of the decentralized inference as we illustrated in Section 5.3.
Appendix F Case Study of the Scheduling Result
We list the partition results generated by HexGen in the heterogeneous-full-price scenario. We use the following representation to describe the scheduled results. We use an array to specify one independent inference pipeline, and the number represents the degree of tensor parallelism for this pipeline stage. For example, [4,2] indicates a two-stage pipeline, where the first stage has a tensor parallel degree of 4, and the second stage has a tensor parallel degree of 2. The scheduled results are listed in Table 4.
Partition breakdown. In Iceland, two 8 $\times$ 3090Ti instances deploy [4,4] strategy to support two model replicas, each function within a single machine. Two 3 $\times$ 3090Ti instances in Norway employ a [2,1,1,2] strategy; cross-machine pipeline parallelism communication happens between the 2nd and 3rd stage. In Nevada, 8 $\times$ A5000 GPUs deploy a [4,4] strategy. In Illinois, 12 $\times$ A6000 GPUs serve four model replicas, each with strategy [2,1]. The remaining 4 $\times$ A6000 / 4 $\times$ A40 is split into four groups, each 2 $\times$ A6000 / 2 $\times$ A40 and another 2 $\times$ A5000 deploys a [2,2] strategy, inter-machine communication is finished by pipeline parallelism.
Table 4: GPU Deployment and Strategy by Region.
Region GPU Configuration Strategy
Iceland $8\times 3090\text{Ti}$ $[4,4]$
$8\times 3090\text{Ti}$ $[4,4]$
Norway $3\times 3090\text{Ti}+3\times 3090\text{Ti}$ $[2,1,1,2]$
[…]
Interesting insight. In a homogeneous setting, the 16 A100 GPUs can serve 4 Llama-2 (70B) model replicas. While in a heterogeneous setting, the 58 cloud GPUs with various types can serve a maximum of 12 Llama-2 (70B) model replicas with various hybrid parallel configurations within the same budget. Here are some interesting insights:
  • In this scenario, although individual inference tasks in a heterogeneous environment may experience increased latency, the overall performance of the system significantly improves.
  • Comprehensively, our scheduling approach prioritizes intra-machine tensor model parallelism to minimize single request latency and employs inter-machine pipeline parallelism to reduce communication over limited bandwidth. It avoids cross-region communication due to ultra-low bandwidth and aims to maximize device memory utilization by incorporating as many model replicas as possible.
  • Additionally, asymmetric parallelism plays a significant role in enhancing system performance, primarily by allowing the adoption of more adaptable parallel strategies to minimize extensive communication via low-bandwidth links. For instance, consider a scenario with 4 A5000 GPUs on one machine and 2 on another. Given the exceptionally low intercommunication bandwidth and the requirement for at least 6×24 GB GPUs to support a 70B model replica, the optimal configuration involves establishing a tensor model parallel group of 4 as the first pipeline stage and a group of 2 as the second. This setup is preferred because pipeline parallel communication demands are substantially lower than those of tensor model parallelism.